# New insight into the orbital parameters of the gamma-ray binary HESS J0632+057

### v1

@@ 3.2.2 Orbital solution with SALT and M18 data
While the new solution is similar to that found from the M18 data alone, it is important to note that with the new orbital period, there is sparse coverage and a larger scatter between phases , which is not covered by the new SALT observations. Periastron lies between these phases. This means that the key parameters of the phase of periastron, the longitude of periastron, and the eccentricity of the orbit are still very poorly constrained (only the statistical errors from the fit are given in Table 2). Increased phase coverage over these critical phases, could significantly improve the orbital solution.
@@ 3.2.3 Orbital solution with SALT and C12 H data
Given the sparse coverage and large scatter in the M18 data around the expected phases of periastron, we have also extracted the RVs reported in fig. 4 of C12, which were obtained from their Liverpool Telescope (LT) spectra of H . However, we note the large offset between the LT RVs and the spectra obtained with the William Herschel Telescope, Mercator Telescope and STELLA-I Telescope also shown in fig. 4 of C12, as well as with respect to SALT. This is likely due to the re-binning that was performed on the LT spectra by C12. In order to remove the systematic offset, we have subtracted the average off-set between the C12 and SALT data - weighted by the scatter of the points in phase bins of width 0.1. In addition, given that the uncertainties in the RV measurements are smaller than the data points, we have estimated the uncertainties to be equivalent to the size of the data points. We then obtained a further orbital solution by combining these data with our average SALT RVs as above. The orbital parameters found are given in Table 2, the best RV fit is shown in the bottom panel of Fig. 4, and the orbit is shown in Fig. 5(b).
@@ 3.4 Implications for the multi-wavelength emission
The discrepancy between the C12 and M18 solutions, has led to difficulty in interpreting the behaviour of HESS J0632+057. The new RVs reported here, along with the new orbital solutions determined with the updated orbital period and the H RV data from M18 and C12, respectively, strongly suggest periastron lies between orbital phases , in contrast to the C12 solution. The orbital solution from the combined SALT and M18 data would place the first peak in the X-ray/TeV light curves (), which is brighter and narrower, after apastron and the broader, fainter second X-ray/Tev peak, at , after periastron.
However, the alternative solution, obtained with the SALT + C12 H data, places the first peak closer to periastron, and the second peak closer to apastron (for a much lower eccentricity). Interestingly, this solution would be in line with that suggested by Malyshev et al. (2019), though the longitude of periastron is different. From X-ray observations, the authors showed that the hydrogen column density, , produces a larger, primary peak at and a smaller, secondary peak at (see Malyshev et al., 2019, and references therein). Since the density of the circumstellar disc decreases with distance from the star, while the scale height increases, the second orbital solution we find would be consistent with this result for an inclined disc, if the emission region is close to the compact object. In this scenario, the peaks in the X-ray/Tev light-curves are produced during the pulsar-disc crossing, as has been previously suggested. The variation in the strength of the emission lines indicates orbit-to-orbit variability of the circumstellar disc. A change in the disc density would change the interaction between the compact object and the circumstellar disc and lead to the orbit-to-orbit variability observed within the non-thermal emission (e.g. Adams et al., 2021).
@@ 4 Conclusions
We present new, independent RV measurements obtained with SALT for the gamma-ray binary HESS J0632+057 where two different, inconsistent orbital solutions have previously been found. The RVs are calculated from Voigt profiles fit to the Balmer emission line wings, and from cross-correlation of filtered regions of the spectra which contained narrow absorption features and low-intensity Fe ii emission lines. A search for RVs from the absorption lines alone did not find a consistent result.
The SALT data are more consistent with the M18 solution. From the SALT data alone, we can constrain the period of periastron to lie between the phases of 0.3-0.6, which disagrees with the phase of periastron obtained in the C12 solution. We present new orbital solutions from the combination of the new SALT observations and the M18 data, which finds a similar solution to M18, but is poorly constrained around periastron. However, combining the SALT observations with the RVs reported for the H emission by C12 we find a less eccentric orbital solution which would provide a better interpretation of the X-ray and TeV light-curves as the pulsar crossing an inclined disc.
@@
Additional observations around phases , not covered by our SALT data, can significantly improve the current orbital solutions.
@@ Acknowledgements
The authors are grateful to P.A. Charles for the very valuable discussions. All of the observations reported in this paper were obtained with the Southern African Large Telescope (SALT) under programme 2020-2-MLT-007 (PI: B. van Soelen). This work was supported by the Department of Science and Technology and the National Research Foundation of South Africa through a block grant to the South African Gamma-Ray Astronomy Consortium. BvS acknowledges support by the National Research Foundation of South Africa (grant number 119430). This work made use of Astropy:111http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration et al., 2013b, 2018, 2022).

### v2

@@ 3.2.2 Orbital solution with SALT and M18 data
While the new solution is similar to that found from the M18 data alone, it is important to note that with the new orbital period, there is sparse coverage and a larger scatter between phases –, which is not covered by the new SALT observations. Periastron lies between these phases. This means that the key parameters of the phase of periastron, the longitude of periastron, and the eccentricity of the orbit are still very poorly constrained (only the statistical errors from the fit are given in Table 2). Increased phase coverage over these critical phases, could significantly improve the orbital solution.
@@ 3.2.3 Orbital solution with SALT and C12 H data
Given the sparse coverage and large scatter in the M18 data around the expected phases of periastron, we have also extracted the RVs reported in fig. 4 of C12, which were obtained from their Liverpool Telescope (LT) spectra of H . However, we note the large offset between the LT RVs and the spectra obtained with the William Herschel Telescope, Mercator Telescope and STELLA-I Telescope also shown in fig. 4 of C12, as well as with respect to SALT. This is likely due to the re-binning that was performed on the LT spectra by C12. In order to remove the systematic offset, we have subtracted the average off-set between the C12 and SALT data – weighted by the scatter of the points in phase bins of width 0.1. In addition, given that the uncertainties in the RV measurements are smaller than the data points, we have estimated the uncertainties to be equivalent to the size of the data points. We then obtained a further orbital solution by combining these data with our average SALT RVs as above. The orbital parameters found are given in Table 2, the best RV fit is shown in the bottom panel of Fig. 4, and the orbit is shown in Fig. 5(b).
@@ 3.4 Implications for the multi-wavelength emission
The discrepancy between the C12 and M18 solutions has led to difficulty in interpreting the behaviour of HESS J0632+057. The new RVs reported here, along with the new orbital solutions determined with the updated orbital period and the H RV data from M18 and C12, respectively, strongly suggest periastron lies between orbital phases –, in contrast to the C12 solution. The orbital solution from the combined SALT and M18 data would place the first peak in the X-ray/TeV light curves (), which is brighter and narrower, after apastron and the broader, fainter second X-ray/Tev peak, at –, after periastron.
However, the alternative solution, obtained with the SALT + C12 H data, places the first peak closer to periastron, and the second peak closer to apastron (for a much lower eccentricity). Interestingly, this solution would be in line with that suggested by Malyshev et al. (2019), though the longitude of periastron is different. From X-ray observations, the authors showed that the hydrogen column density, , produces a larger, primary peak at and a smaller, secondary peak at (see Malyshev et al., 2019, and references therein). Since the density of the circumstellar disc decreases with distance from the star, while the scale height increases, the second orbital solution we find would be consistent with this result for an inclined disc, if the emission region is close to the compact object. In this scenario, the peaks in the X-ray/Tev light curves are produced during the pulsar-disc crossing, as has been previously suggested. The variation in the strength of the emission lines indicates orbit-to-orbit variability of the circumstellar disc. A change in the disc density would change the interaction between the compact object and the circumstellar disc and lead to the orbit-to-orbit variability observed within the non-thermal emission (e.g. Adams et al., 2021).
@@ 4 Conclusions
We present new, independent RV measurements obtained with SALT for the gamma-ray binary HESS J0632+057 where two different, inconsistent orbital solutions have previously been found. The RVs are calculated from Voigt profiles fit to the Balmer emission-line wings, and from cross-correlation of filtered regions of the spectra which contained narrow absorption features and low-intensity Fe ii emission lines. A search for RVs from the absorption lines alone did not find a consistent result.
The SALT data are more consistent with the M18 solution. From the SALT data alone, we can constrain the period of periastron to lie between the phases of 0.3-0.6, which disagrees with the phase of periastron obtained in the C12 solution. We present new orbital solutions from the combination of the new SALT observations and the M18 data, which finds a similar solution to M18, but is poorly constrained around periastron. However, combining the SALT observations with the RVs reported for the H emission by C12 we find a less eccentric orbital solution which would provide a better interpretation of the X-ray and TeV light curves as the pulsar crossing an inclined disc.
@@
Additional observations around phases –, not covered by our SALT data, can significantly improve the current orbital solutions.
@@ Acknowledgements
The authors are grateful to PA Charles for the very valuable discussions. All of the observations reported in this paper were obtained with the SALT under programme 2020-2-MLT-007 (PI: B. van Soelen). This work was supported by the Department of Science and Technology and the National Research Foundation of South Africa through a block grant to the South African Gamma-Ray Astronomy Consortium. BvS acknowledges support by the National Research Foundation of South Africa (grant number 119430). This work made use of astropy:111http://www.astropy.org a community-developed core python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration et al., 2013, 2018, 2022).